### v1

# Superpixel Tokenization for Vision Transformers: Preserving Semantic Integrity in Visual Tokens

Abstract
Transformers, a groundbreaking architecture proposed for Natural Language Processing (NLP), have also achieved remarkable success in Computer Vision. A cornerstone of their success lies in the attention mechanism, which models relationships among tokens. While the tokenization process in NLP inherently ensures that a single token does not contain multiple semantics, the tokenization of Vision Transformer (ViT) utilizes tokens from uniformly partitioned square image patches, which may result in an arbitrary mixing of visual concepts in a token. In this work, we propose to substitute the grid-based tokenization in ViT with superpixel tokenization, which employs superpixels to generate a token that encapsulates a sole visual concept. Unfortunately, the diverse shapes, sizes, and locations of superpixels make integrating superpixels into ViT tokenization rather challenging. Our tokenization pipeline, comprised of pre-aggregate extraction and superpixel-aware aggregation, overcomes the challenges that arise in superpixel tokenization. Extensive experiments demonstrate that our approach, which exhibits strong compatibility with existing frameworks, enhances the accuracy and robustness of ViT on various downstream tasks. The source codes are available at: https://github.com/jangsoohyuk/SuiT
1 Introduction
Figure 1: A high-level overview of tokenization. (Top) The conventional grid-like tokenization in ViT and (Bottom) our proposed superpixel tokenization.
Transformers [51], originally proposed for Natural Language Processing (NLP), have demonstrated exceptional performance across diverse tasks such as machine translation [55, 19, 5], natural language understanding [27, 34], and text generation [41, 31]. The transformative power of Transformers was again attested by the release of large language models [7], which continue to revolutionize NLP as we know it. Similarly, the Vision Transformer (ViT) [14], which extended the Transformer architecture to image recognition, changed the landscape of Computer Vision research. The subsequent research efforts to broaden the usage of ViTs to various visual tasks, including detection [45, 8, 65], segmentation [32, 63, 48], generation [15, 38, 60], and multi-modal applications [42], have made the ViT the dominant architecture in Computer Vision.
In NLP, the success of Transformers is commonly attributed to the attention mechanism, which captures contextual dependencies by modeling the relationship among input texts, which are expressed as tokens. Widely adopted tokenization methods in NLP, such as Byte Pair Encoding [46], WordPiece [47], and SentencePiece [29], first split each sentence into words, which are then broken down into more granular tokens. For instance, when tokenizing “Attention is All You Need,” each word is separated into: [“Attention,” “is,” “All,” “You,” “Need”]. By definition, this initial word-level splitting of sentences prevents multiple words with different meanings from being fused into a single token.
In contrast, fixed-size grid patches form the basis of tokenization in ViT and its variants [14, 50, 35]. As visualized in Figure 1, when naïvely sliced square image patches are used as tokens, a single token can simultaneously contain two or more visual concepts, e.g., eye and fur. Consequently, this approach in ViT diverges from the core principle of tokenization, which is to convert plain inputs into smaller yet more structured subunits for further processing.
In this work, we employ the concept of superpixels to design a tokenization process that yields semantically preserving tokens that contain a single, isolated visual concept. A superpixel is a cluster of connected pixels in an image that share similar properties, such as color, texture, or position, and it is often used as a fundamental unit in image processing [2, 11, 33]. Figure 1 illustrates how leveraging superpixels alters tokenization of ViT. Grouping pixels into superpixels divides an image into cohesive regions that capture similar semantics, thereby preventing the mixture of unrelated semantic information in each token.
Despite the advantage of superpixels in tokenization, the different shapes and sizes of each superpixel present significant challenges in utilizing them as ViT tokens [25, 59, 62]. In ViTs, tokens are generated from fixed-size square patches at predefined locations through simple flattening and linear projection. Because superpixels consist of varying numbers of pixels, de-facto flattening and linear projection operations in ViT are incompatible with superpixels. Moreover, the location of each superpixel changes dynamically depending on the image. The variability of superpixel locations makes it infeasible to directly apply the positional embeddings of ViT, which are designed for fixed locations. The inadequacy of original ViT operations necessitates novel techniques for addressing irregular geometry and composition to reap the inherent benefits of superpixels.
We thus propose a novel tokenization pipeline that overcomes the aforementioned problems and enables the effective incorporation of superpixels into ViT tokenization. The proposed pipeline consists of two major technical components: pre-aggregate feature extraction and superpixel-aware feature aggregation. The first stage prepares pre-aggregate features for the subsequent superpixel-aware aggregation, which utilizes pooling operations to remove the irregularity and variability of superpixels. Applying pooling operations directly on the input image may result in a severe loss of information. By training the pre-aggregate feature extractor and applying it prior to our feature aggregation, we effectively sidestep this potential information loss. Collectively, this two-stage process allows the proposed tokenization approach to leverage the benefits of superpixels while preserving important details of an image.
We demonstrate the superiority of our Superpixel-Tokenized Vision Transformer (SuiT) compared to the baselines. SuiT outperforms in a variety of tasks including ImageNet-1K classification [12], segmentation [56], transfer learning [66], and self-supervised learning [9]. We further analyze the properties that emerge using our tokenization pipeline by conducting extra examinations. Furthermore, we visualize that our proposed tokenization method indeed results in image tokens that preserve semantics more effectively than the grid-based tokenization method.
Our contribution can be summarized as follows:
  • We propose an effective superpixel-based tokenization method that overcomes the challenges associated with the use of superpixels.
  • We empirically show that our tokenization method produces tokens that better preserve semantic information than grid-based tokens.
  • We demonstrate the effectiveness of our approach through experiments on image classification, transfer learning, self-supervised learning, and zero-shot segmentation.
2 Related Work
Superpixel-based Vision Models
Superpixels, which are perceptually meaningful clusters of pixels, have been a foundational concept in computer vision [2, 11, 33]. Previous studies have attempted to use superpixels in several tasks such as segmentation [30, 37] and object detection [18], but these works did not demonstrate the broad applicability of superpixels across a range of vision tasks.
Another recent work, STViT [23], applies superpixel-like clustering across all layers via supertoken attention, but operates at the token level, resulting in a coarser granularity than superpixels. CoC [36], borrows the concept of superpixels within neural networks but lacks further exploration of superpixel tokenization. Similarly, SPFormer [37] utilizes superpixel-based representations but requires specialized attention modules making it difficult to integrate with existing ViT architectures in a plug-and-play manner.
In contrast to these approaches, our research explores a simple yet effective superpixel tokenization method that can be easily employed in a transformer architecture for a range of general-purpose tasks.
Tokenization Methods
While most ViT variants focus on improving backbone architectures and attention mechanisms using naive square-shaped patches, there have been several attempts in exploration of advanced tokenization strategies tailored for vision tasks. Quadformer [44] and MSViT [16] both propose adaptive tokenization strategies that dynamically adjust token resolutions based on image content.
The studies SPiT [1] and sViT [28] are the most similar to our work in that they target only the tokenizing module while keeping the backbone intact and share similar research motivations. SPiT [1] focused on analyzing this tokenization from the explainable AI (XAI) perspective but did not explore its generalizability across various tasks. Also, SPiT [1] suffers from performance degradation without gradient features, as they lack of abundant features due to their learning free feature extraction. Moreover, its applicability to other learning paradigms, such as self-supervised learning [9], was not investigated.
sViT [28] shared a similar motivation in aiming for semantically preserving tokenization, but it still divides tokens based on a naïve bounding box, failing to create truly semantically preserving tokens. Additionally, resizing these tokens into square shapes compromised the visual structure of objects within the tokens. These shortcomings in design leads leads to limited performance in various tasks.
As described in the previous section, we propose a semantically preserving tokenization through superpixels in vision and investigate the performance of our method across various tasks and explore its adaptability to self-supervised frameworks.
Figure 2: Overview of our superpixel tokenization pipeline. Local features are extracted and combined with positional encodings, followed by superpixel-aware aggregation using mean and max pooling to produce superpixel tokens, which are fed into a Vision Transformer.
3 Preliminaries
In this section, we introduce the key concepts and notations that are integral to our work, which aims to construct semantically preserving ViT tokens with superpixels. In Section 3.1, we describe the mechanism of superpixel algorithms, and in Section 3.2, we go over the traditional ViT tokenization. Subsequently, we discuss the research challenges in adopting superpixels as tokens in ViTs.
3.1 Superpixel Algorithms
Superpixels are clusters of image pixels grouped together based on color, position, or other similarity measures. They provide a compact and meaningful representation of the image by grouping pixels into regions that align with its underlying structure.
Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, a superpixel algorithm $\mathcal{S}$ produces a superpixel index map $\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$:
$\mathcal{I}=\mathcal{S}(\mathbf{x},K,m),$ (1)
where $K$ is the number of superpixels, and $m$ is a parameter that controls the trade-off between color and positional distance, i.e., compactness in SLIC [2].
The superpixel index of at coordinate $(h,w)$ can be defined as $\mathcal{I}^{(h,w)}$, where $0<h\leq H,0<w\leq W$. A $k$-th superpixel $\mathcal{C}_{k}$ can then be defined as:
$\mathcal{C}_{k}=\{\textbf{x}^{(h,w)}\mid\mathcal{I}^{(h,w)}=k,0<k\leq K\},$ (2)
where $\textbf{x}^{(h,w)}$ is the pixel at $(h,w)$. Each superpixel $\mathcal{C}_{k}$ consists of varying numbers of pixels with different shapes and sizes.
3.2 Tokenization in Vision Transformers
In grid-based tokenization of conventional ViT and its variants [50, 35, 42], the input image $\mathbf{x}$ is divided into $N=\frac{H\times W}{P^{2}}$ non-overlapping patches of size $P\times P$, each represented as $\mathbf{x}^{(i)}\in\mathbb{R}^{P\times P\times C}$, where $\mathit{i}\in\{1,...,N\}$ corresponds to the index of each patch. Each patch $\mathbf{x}^{(i)}$ is flattened into a vector $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{P^{2}C}$ and linearly projected using an embedding matrix $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$, where $D$ is the dimension of tokens. In addition to this, positional embedding $\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$ of its location is injected to form the input token $\mathbf{z}^{(i)}=\mathbf{x}_{p}^{(i)}\mathbf{E}+\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$. The set of tokens $\mathbf{Z}=\{\mathbf{z}^{(i)}\mid i\in\{1,2,...,N\}\}$ which originates from patches as above, is given to the ViT models along with a learnable classification token $\mathbf{z}_{\text{cls}}$ for classification.
Challenges
The tokenization process in ViT requires square-shaped and fixed-sized patches, acquired at uniform locations. These constraints on image patches make a naïve application of superpixels to ViT tokenization challenging. First, the straightforward flatten-and-project operation is incompatible with superpixels that come in varying sizes and shapes [25, 59, 62]. Flattening of superpixels with diverse shapes and varying numbers of pixels yields vectors of varying lengths, which hinder a simple linear projection. Second, unlike grid patches with fixed positions, superpixels are located at different positions depending on the image. Consequently, the standard positional embedding, which relies on fixed positions, is inappropriate for superpixels.
4 Method
To address the aforementioned issues in Section 3, we propose a novel superpixel tokenization pipeline that effectively embeds each superpixel into a single token. The overview of superpixel tokenization is illustrated in Figure 2. Our proposed pipeline incorporates two components: pre-aggregate feature extraction (Section 4.1) and superpixel-aware feature aggregation (Section 4.2). Our aggregation is based on pooling operations to address the inconsistent size and shape of superpixels. However, naïve application of such aggregation on the RGB space may lead to critical loss of information. To sidestep this potential risk, we train a pre-aggregate feature extractor and use its output as input for the superpixel-aware aggregation. These two processes of the proposed tokenization method together exploit the benefits of superpixel-based tokenization while conserving critical image information.
4.1 Pre-aggregate Feature Extraction
Local Feature Extraction
We start by extracting local features from the input image using a simple convolutional block. We adopt the design of the initial convolutional block in ResNet [17], which consists of a $7\times 7$ convolutional layer, a Batch Normalization layer [24], and a ReLU activation function [3]. We make a slight modification by replacing ReLU with GELU [20], to align with recent advances in the field [14, 35]. Given an input image $\mathbf{x}$ and the convolutional block, we extract local features $\mathbf{F}_{local}\in\mathbb{R}^{H\times W\times d_{local}}$.
Positional Features
Positional features play a crucial role in transformers, as they are unaware of the order of input tokens. While the original Transformer proposed to use sinusoidal positional encodings with pre-defined frequencies [51], ViTs opted to use learnable positional embeddings for each pre-defined locations [14, 50].
Superpixels have complex positional information of high granularity. Using the learnable positional embeddings as in ViT is inadequate, as it would require excessive number of parameters, i.e., $H\times W$ learnable embeddings each of $D$-dimensions. To keep it parameter-efficient, we adopt the sinusoidal positional encoding with learnable frequencies, which is known to capture high frequency information efficiently [49]. To our knowledge, this positional encoding scheme has not been applied to the Transformer literature before, and is a peculiar choice to adapt to our design of using superpixels of high-granularity.
Consequently, we acquire the value of $q$-th dimension in $d_{pos}$-dimensional positional feature $\mathbf{F}_{pos}\in\mathbb{R}^{H\times W\times d_{pos}}$ at each spatial coordinate $(h,w)$ as below:
$\begin{split}&\mathbf{F}^{(h,w)}_{pos}[2q]=\sin(f_{y}[q]\cdot h+f_{x}[q]\cdot w),\\ &\mathbf{F}^{(h,w)}_{pos}[2q+1]=\cos(f_{y}[q]\cdot h+f_{x}[q]\cdot w),\end{split}$ (3)
where $f_{x}[q]$ and $f_{y}[q]$ each denote the frequency of horizontal and vertical axis in the $q$-th dimension.
Projection
Given the local features $\mathbf{F}_{local}$ and the positional features $\mathbf{F}_{pos}$, we integrate the two to obtain our pre-aggregate features. The two feature maps $\mathbf{F}_{local}$ and $\mathbf{F}_{pos}$ are concatenated and passed to a linear projection layer $\mathcal{P}$:
$\mathbf{F}=\mathcal{P}(\langle\mathbf{F}_{local},\mathbf{F}_{pos}\rangle),$ (4)
[…]
4.2 Superpixel-aware Feature Aggregation
To aggregate the extracted pre-aggregate features within each superpixel into a representative embedding, we employ average and max pooling, which effectively handles features of varying numbers. The pre-computed superpixel index map $\mathcal{I}$ of an input image x is used as the guideline for determining which features belong to which superpixels.
Given the $D/2$-dimensional pre-aggregate features, we perform average pooling and max pooling within each superpixel cluster. This results in two feature vectors per superpixel: average features $\mathbf{z}_{\text{avg}}$ and max features $\mathbf{z}_{\text{max}}$, each of $D/2$-dimensions:
$\begin{split}\mathbf{z}^{(k)}_{\text{avg}}&=\frac{1}{|\mathcal{C}_{k}|}\sum_{\textbf{x}^{(h,w)}\in\mathcal{C}_{k}}\mathbf{F}^{(h,w)},\\ \mathbf{z}^{(k)}_{\text{max}}&=\max_{\textbf{x}^{(h,w)}\in\mathcal{C}_{k}}\mathbf{F}^{(h,w)},\end{split}$ (5)
where $|\mathcal{C}_{k}|$ and $\mathbf{F}^{(h,w)}$ denote the number of pixels within superpixel $\mathcal{C}_{k}$ and the pre-aggregate feature at coordinate $(h,w)$, respectively.
Average pooling and max pooling extract complementary features from each superpixel. Average pooling captures the overall characteristics common in the superpixel, while max pooling identifies the most salient features that may be obscured by averaging alone. We concatenate $D/2$-dimensional vectors from each pooling method along the channel dimension, yielding a $D$-dimensional token $\mathbf{z}^{(k)}$:
$\mathbf{z}^{(k)}=\langle\mathbf{z}^{(k)}_{\text{avg}},\mathbf{z}^{(k)}_{\text{max}}\rangle.$ (6)
This approach effectively combines both general and prominent information, creating a more comprehensive representation of each superpixel.
5 Experiments
Size Model # Params. GMACs # tokens Top-1 Tiny DeiT 5.7M 1.26 196 72.2 SuiT 5.5M 0.79 100 72.2 1.26 171 75.3 1.44 196 75.7 Small DeiT 22.1M 4.61 196 79.8 SuiT 21.7M 3.64 132 79.8 4.60 172 80.5 5.20 196 80.9 Base DeiT 86.8M 17.6 196 81.8 SuiT 86.0M 16.4 160 81.8 17.6 173 82.0 19.7 196 82.1
Table 1: Comparison of classification performance on ImageNet-1K against DeiT. The performance of SuiT is reported with varying token counts. Underlined values indicate those that match DeiT’s GMACs or token counts.
5.1 Experimental Settings
We validate our tokenization method by applying it to ViTs, dubbed as SuiT. We experiment on three model scales, Tiny ($D=192$), Small ($D=384$), and Base ($D=768$). The values $d_{local}$ and $d_{pos}$ are both set to $D/4$.
We employ FastSLIC [4] as our superpixel algorithm due to its high speed. We use a stride of 2 in the convolutional layer of local feature extraction, which produces feature maps of 2×2$\times$2 × smaller size than the input. The positional feature map $\mathbf{F}_{pos}$ and superpixel index map $\mathcal{I}$ are scaled to the same size accordingly.
This design choice was based on the observation that using a stride of 1 offered no notable performance improvements while significantly increasing computational costs. Further details are described in Appendix.
5.2 Comparison to Baseline
As our approach focuses on tokenization without modifying the backbone network, we focus on comparison with our baseline, DeiT [50] on ImageNet-1K classification [12]. We mostly follow the experimental settings from DeiT, with minimal changes for fair comparison.
Table 1 compares the classification performance of SuiT to that of DeiT. For each model size, the table provides a detailed comparison across key aspects: classification accuracy, computational cost measured in GMACs, and the number of tokens processed. This adaptive inference capability arises from the flexibility to adjust the number of superpixel tokens, enabling users to customize the configuration based on specific needs. When achieving comparable performance, SuiT exhibits substantially higher efficiency than DeiT. At equivalent GMACs, SuiT delivers superior accuracy. Maintaining an identical number of tokens further boosts its accuracy,
Tiny Small Base
Quadformer [44] - 80.8 84.4
MSViT [16] 75.5 82.0 -
SuiT (Fine-tuned) 76.1 82.6 84.5
SPiT [1] - 75.0 80.4
SuiT (Scratch) 75.7 80.9 82.1
Table 2: Comparison with existing studies on tokenization. Experiment above reports the Top-1 accuracy on ImageNet-1K. The numbers for Quadformer-Tiny, MSViT-Base and SPiT-tiny are intentionally left blank, as they are unavailable.
Size Model $\text{INat}_{18}$ $\text{INat}_{19}$ Cars Flowers
Small DeiT 70.7 76.6 92.1 96.5
SuiT 71.4 77.9 92.1 96.8
Base DeiT 73.2 77.7 92.1 96.8
SuiT 74.0 78.4 92.1 96.9
DINO-ViT 72.6 78.6 93.0 97.1
DINO-SuiT 72.8 79.4 93.8 97.2
Table 3: Results of transfer learning different datasets in terms of top-1 accuracy results. SuiT outperforms ViT in most downstream tasks in both supervised and self-supervised settings. All results for DeiT and DINO are reproduced following the training recipes provided by the authors. : The reported Flowers102 performance could not be reproduced in the exact setting, so we compare with our reproduced results.
ECSSD DUTS DUT-OMRON
Model Postproc. IoU Acc. max $F_{\beta}$ IoU Acc. max $F_{\beta}$ IoU Acc. max $F_{\beta}$
SPiT [1] $\times$ 77.3 93.4 90.3 63.9 89.4 77.1 56.4 86.8 71.1
DINO-ViT [9] 76.7 93.2 86.9 61.0 90.6 74.1 59.9 88.5 67.6
DINO-ViT [9] $\times$ 71.1 91.6 80.4 56.7 89.5 66.4 51.8 86.8 58.5
DINO-SuiT 79.7 93.8 88.7 60.2 89.1 71.3 60.5 87.9 67.6
DINO-SuiT $\times$ 80.5 ($\uparrow$ 9.4) 93.8 ($\uparrow$ 2.2) 87.0 ($\uparrow$ 6.6) 60.0 ($\uparrow$ 3.3) 88.8 ($\downarrow$ 0.7) 68.0 ($\uparrow$ 1.6) 57.5 ($\uparrow$ 5.7) 87.2 ($\uparrow$ 0.4) 63.4 ($\uparrow$ 4.9)
Table 4: Results of zero-shot salient segmentation with TokenCut [56]. Models with additional postprocessing are colored in gray. The best performance in each column is highlighted in bold, and the second-best performance is indicated with an underline among non-postprocessed results. denotes that the model is trained at a different training setting to others, i.e., with different resolution ($256$ vs $224$), data augmentation techniques and learning method (supervised vs self-supervised).
5.3 Comparison to Existing Tokenization Methods
We evaluate our tokenization method against existing approaches under two scenarios used in existing tokenization studies. For a fair comparison, we follow the setup of  [44, 16], which fine-tunes a ViT pretrained on ImageNet-21K [43]. Additionally, we include SPiT  [1], a concurrent work leveraging superpixels for tokenization, which trains its model from scratch on ImageNet-1K. Both experiments are conducted on ImageNet-1K classification tasks.
As reported in Table 2, our method consistently outperforms existing tokenization methods across all scales of model size, showing its efficacy. On fine-tuning experiments following [44, 16], SuiT outperforms existing tokenization methods for fine-grained tokens, such as Quadformer [44] and MSViT [16].
In experiments for training from scratch following [1], SuiT consistently outperforms SPiT, another work on superpixel-based tokenization, by a noticeable margin. These results verify the excellence of our tokenization pipeline compared to prior work.
5.4 Transfer Learning
We verify the generalization ability of SuiT when transferred to various downstream tasks. We follow the standard fine-tuning protocol and default settings of DeiT [50] for fair comparison, by fine-tuning ImageNet-1K pre-trained SuiT to each downstream task specifically. We test our models on iNaturalist [22], Flowers102 [39], and StanfordCars [10]. Further details can be found in the Appendix.
Table 3 presents a comparison of the classification accuracy between DeiT and SuiT when fine-tuned on various downstream tasks. Across various tasks and model scales,  SuiT consistently achieves higher accuracy than DeiT. Notably, on iNaturalist2018, SuiT-Small outperforms DeiT-Small by 0.8%p, demonstrating its superior ability to effectively transfer learned features, particularly when ample data is available.
Figure 3: Self-attention map comparison between DINO-ViT and DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
Figure 4: Qualitative results of zero-shot salient object segmentation. denotes the model with additional post-processing. DINO-SuiT successfully detects salient objects both in single- and multi-object scenarios without any post-processing.
These results indicate that SuiT shows improved generalizability than DeiT. This positions SuiT a strong choice for transfer learning across diverse domains. The consistently superior performance of SuiT, regardless of model size or dataset complexity, further underscores its suitability for real-world applications.
5.5 Self-supervised Learning
We further study the applicability of SuiT in self-supervised scenario. Specifically, we train SuiT on ImageNet-1K with DINO [9], a well-known self-supervised learning approach, chosen for its proven effectiveness and popularity.
5.5.1 Transfer Learning
After pre-training SuiT with DINO, we evaluate its performance when transferred to downstream tasks to confirm the quality of learned features. As shown in the lower section of Table 3, DINO-SuiT consistently surpasses DINO-ViT across all downstream tasks. This demonstrates that SuiT effectively supports self-supervised scenarios, further establishing its strength as a versatile pipeline for a wide range of vision tasks.
5.5.2 Zero-shot Salient Object Segmentation
It is well known that the attention map of ViTs trained with DINO effectively attends to object regions without explicit supervision [9]. As shown in Figure 3, while DINO-ViT’s self-attention maps are limited to square patch-level details, those of DINO-SuiT even capture finer structures of salient objects. To demonstrate that this fine-grained property synergistically enhances the salient object segmentation capabilities of DINO [9], we measure the segmentation performance on three benchmark datasets, ECSSD [57], DUTS [54] and DUT-OMRON [58] following the works of Wang et al. [56], Aasan et al. [1]. We conduct our experiments on the TokenCut framework [56], which is a normalized graph cut algorithm that extracts saliency mask from attention map.
Superpixel tokenization significantly improves DINO-ViT [9] performance across most metrics without post-processing, achieving results comparable to DINO-ViT with post-processing.
Notably, as shown in Table 4, our model exhibits minimal performance differences between post-processed and non-post-processed cases, unlike DINO-ViT. Compared to SPiT [1], SuiT demonstrates competitive performance across three datasets. However, SPiT [1] does not show its applicability to self-supervised methods such as DINO and is trained with larger resolution images. Qualitative comparisons in Figure 4 highlight that DINO-SuiT detects all salient objects, even in multi-object scenarios, whereas DINO-ViT shows limitations such as rectangular masks without post-processing and incomplete object detection.
Figure 5: Comparison of K-means clustering results for patch tokens from DeiT and superpixel tokens from SuiT, both supervisedly trained on ImageNet-1K, with K=6. Superpixel tokens in SuiT tend to cluster by semantic meaning, while ViT patch tokens mostly cluster by positional information.
6 Analysis
6.1 Semantic Integrity of Superpixel Tokens
To verify the improved semantic preservation of our superpixel tokens, we perform K-means clustering [26] on tokens embedding space. We compare the clustering results of superpixel tokens with grid-based tokens.
As shown in Figure 5, our superpixel tokens are grouped based on semantic similarity rather than spatial proximity, unlike the patch tokens from DeiT, which lack consistency within clusters. For example, in the first row, tokens corresponding to the regions of ducklings are grouped together, whereas clusters from DeiT’s patch tokens fail to convey clear meaning. This analysis suggests that superpixel tokens are better at preserving semantics. Additional results are provided in the Appendix.
Figure 6: Class-identifiability analysis results of DeiT and SuiT. We show the proportion of class-identifiable [CLS] tokens and image tokens for SuiT and DeiT across all layers.
Size Model ImageNet-A ImageNet-O
(Acc.) (AUPR)
Tiny DeiT 7.0 17.4
SuiT 11.1 19.8
Small DeiT 19.2 20.9
SuiT 22.5 23.3
Base DeiT 27.9 24.6
SuiT 29.5 26.7
Table 5: Experimental results of SuiT and DeiT on ImageNet-A and ImageNet-O. We report accuracy on ImageNet-A and AUPR on ImageNet-O, following  [21].
6.2 Class Identifiability Analysis
Vilas et al. [52] examined class-specific encoding in supervised ViTs by projecting intermediate representations through the trained classifier head into a class-specific embedding space, similar to the logit lens method in NLP [40, 6]. Building on this, we evaluate class-specific encoding across SuiT’s layers by projecting token representations through the classifier head and measuring the proportion of tokens correctly predicting the target class.
Figure 6 shows that SuiT’s classification token shows higher class identifiability in the earlier layers compared to DeiT. This can be attributed to the semantic preserving nature of superpixel tokenization, which aids in aggregating class-specific encoding in early layers. In the deeper layers, DeiT’s patch tokens exhibit class identifiability levels slightly lower than its classification token, with a marginal gap. This observation aligns with findings from previous studies [52, 53], which report that patch tokens in supervised ViTs tend to encode increasingly global information as layers deepen. In contrast, SuiT’s patch tokens maintain a lower proportion of class-identifiable features in the later layers, suggesting a reduced emphasis on global feature encoding. This divergence alludes that SuiT and DeiT adopt distinctive strategies to aggregate class-specific features at different network depths.
6.3 Enhanced Robustness
We examine the robustness and out-of-distribution (OOD) generalization ability of SuiT in ImageNet variants, ImageNet-A [13] and ImageNet-O [21]. ImageNet-A is a subset of ImageNet-1K curated to mislead classifiers and test model robustness, while ImageNet-O, composed of OOD samples not in ImageNet-1K, evaluates the OOD detection capabilities of models.
In Table 5, SuiT consistently outperforms DeiT on both datasets across model variants.
These findings indicate that our superpixel tokenization improves model robustness in challenging scenarios by mitigating the model’s reliance on features that are easily learned but lack generalizability.
6.4 Ablation Studies
Local feats. PE Freqs. PE Injection Top-1
random init. concat+proj 75.7
random init. concat+proj 66.5
No PE 71.5
predefined concat+proj 73.0
random init. addition 75.2
random init. concat 73.8
Table 6: Ablation study on technical components of pre-aggregate feature extraction.
Avg Std Softmax Max Top-1
75.7
73.2
75.5
Fail
75.5
Table 7: Ablation on our aggregation method. ‘Fail’ indicates that the corresponding approach was unsuccessful in training.
Pre-aggregate Features
The following ablation studies on our proposed components are conducted using SuiT-Tiny. Table 6 highlights the importance of components in pre-aggregate feature extraction. First, using raw 3-channel RGB features leads to a significant performance drop, confirming the need for higher-dimensional local features. Second, removing positional information reduces performance but retains partial class prediction through local cues. Experiments also reveal that predefined positional frequencies under-perform compared to learned frequencies, emphasizing the need for vision-specific customization. Lastly, our results show that the combination of concatenation and projection more effectively integrate local and positional features than simple concatenation or addition used in prior studies [51, 14, 50].
Aggregation
Table 7 presents the analysis of our aggregation method, which employs cardinality-independent operations. Removing max pooling results in a notable performance drop, and removing average pooling causes a slight decrease. We also tested standard deviation as an alternative to max pooling, but it proved numerically unstable and unsuitable for training. Softmax pooling, which computes a weighted sum of values within each superpixel, showed promising results but slightly underperformed compared to max pooling. This may be due to the softmax operation suppressing prominent features crucial for recognition, leading to less discriminative representations.
7 Conclusion
In this work, we proposed a novel superpixel-based tokenization method for ViTs as an attempt to improve the visual tokenization process. By leveraging superpixels, the proposed tokenization obtains semantic-preserving tokens, overcoming the issue of mixed semantics in grid-based patches. To incorporate superpixels that are irregular in shape, size, and location, we proposed a two-stage pipeline consisting of pre-aggregate feature extraction and superpixel-aware feature aggregation. Extensive experiments demonstrated the effectiveness of our method on diverse tasks, e.g., supervised classification, transfer learning, and self-supervised learning. Additional analytical findings provided meaningful insights into the mechanism and emergent properties of superpixel tokenization. We hope our promising results inspire further exploration of token design in ViTs, fostering advancements in both model performance and interpretability.
Supplementary Material
A1 Implementation Details
A1.1 ImageNet-1K classification
A1.1.1 Training from scratch
The settings for experiments from scratch is presented in Table A2. We mostly follow the settings of DeiT [50] with minimal changes. The main difference to the DeiT setting is as follows: 1) to align with our main concept of using superpixels, we disable MixUp [61]. 2) In random erasing [64], the randomly erased region is filled with per-channel random color instead of per-pixel random color. 3) We use gradient clipping, which was known to be helpful in ViT [14], but later removed in DeiT [50]. Other than that, we slightly tweak hyper-parameters of learning rate, weight decay and stochastic depth. For models of all three scales, Tiny, Small and Base, we use a base learning rate of $48e-$ 4, which is scaled proportionally to the batch size, by ×$\times\frac{batchsize}{512}$. We use different weight decay and stochastic depth values for each scale. Most notably, our model uses a larger weight decay and drop rate for stochastic depth, especially in Base scale. This is because we found that our model tends to overfit quickly, compared to the baseline DeiT. We thus opt to use a stronger regularization term. Similar observations were made in Transfer Learning, which is also discussed in the Section  A1.2.
A1.1.2 Fine-tuning from Pre-trained weights
[…]
A1.2 Transfer Learning
Here, we list the training settings used for transfer learning. For both ImageNet-1K [12] classification models and DINO [9] pre-trained models, we follow the default training settings of DeiT [50] and DINO [9]. The details for each experiment can be found at Table A4 and Table A4. Our experiments, especially on smaller datasets, required stronger regularization terms. A very high drop rate for stochastic depth of 0.5 was essential in Flowers and StandfordCars, to avoid overfitting.
Model size Tiny Small Base
Epochs 300
Batch size 1024
Optimizer AdamW
Base Learning rate 8e-4
Learning rate decay cosine
Weight decay 0.02 0.05 0.08
Warmup epochs 5
Label smoothing 0.1
Dropout
Stoch. Depth 0.0 0.1 0.3
Repeated Aug
Gradient Clip.
Rand Augment 9 / 0.5
Mixup prob. 0.0
Cutmix prob. 1.0
Erasing prob. 0.1
Table A1: Settings of SuiT on ImageNet-1K classification, training from scratch. We mostly follow the settings open to public by the authors [50].
Model size Tiny Small Base
Epochs 130
Batch size 1024 1024 400
Optimizer AdamW
Base Learning rate 2e-4 2e-4 5e-5
Learning rate decay cosine
Weight decay 1e-8
Warmup epochs 5
Label smoothing 0.1
Dropout
Stoch. Depth 0.0 0.1 0.3
Repeated Aug
Gradient Clip.
Rand Augment 9 / 0.5
Mixup prob. 0.0
Cutmix prob. 1.0
Erasing prob. 0.1
Table A2: Settings of SuiT on ImageNet-1K classification, fine-tuning from pre-trained weights. We mostly follow the settings open to public by the authors [50].
Datasets iNatuarlist-18, -19 Flowers, StanfordCars
Epochs 300 1000
Batch size 768 768
Optimizer AdamW AdamW
Base learning rate 5e-5 1e-4
Learning rate decay cosine cosine
Weight decay 1e-6 5e-2
Warmup epochs 5 5
Label smoothing 0.1 0.1
Dropout
Stoch. Depth 0.5 0.5
Repeated Aug
Gradient Clip.
Rand Augment 9/0.5 9/0.5
Mixup prob. 0.8 0.8
Cutmix prob. 1.0 1.0
Erasing prob. 0.1 0
Table A3: Training settings for transfer learning ImageNet-1K pre-trained SuiT. We mostly follow the settings open to public by the authors [50].
Datasets iNatuarlist-18, -19 Flowers, StanfordCars
Epochs 300 1000
Batch size 768 768
Optimizer AdamW AdamW
Base learning rate 5e-5 1e-4
Learning rate decay cosine cosine
Weight decay 1e-4 1e-2
Warmup epochs 5 5
Label smoothing 0.1 0.1
Dropout
Stoch. Depth 0.1 0.5
Repeated Aug
Gradient Clip.
Rand Augment 9/0.5 9/0.5
Mixup prob. 0.8 0.8
Cutmix prob. 1.0 1.0
Erasing prob. 0.1 0.25
Table A4: Training settings for transfer learning DINO pre-trained SuiT. We mostly follow the settings open to public by the authors [9].
A1.3 Self-Supervised Learning
For DINO [9] pre-training, we follow the exact training hyper-parameters used to pre-train ViTs, without any tuning, extra training techniques or augmentations. We train DINO-SuiT for 300 epochs, and transfer the pre-trained model to downstream tasks following the settings mentioned in Section A1.2. Additional self-attention maps of DINO-SuiT can be found in Figure A1.
A1.4 Zero-shot Salient Object Segmentation
In the TokenCut [56] algorithm, we searched for and used the best $\tau$ value for both DINO-ViT and DINO-SuiT. For DINO-ViT, $\tau=0.2$ was used, while for DINO-SuiT, $\tau=0.4$ was used. All other settings followed those described in the TokenCut [56] paper. More qualitative results are shown in Figure A2.
A2 Semantic Integrity of Superpixel Tokens
In our semantic integrity analysis, we performed $k$-means clustering on each token embeddings from DeiT [50] and SuiT. In Figure A3, we visualized how tokens were clustered for each method by varying the value of $k$ from 2 to 10. This further shows that our superpixel tokens better preserve the semantics compared to that of DeiT. Most notably, when $k=2$, the tokens are clustered to two groups, foreground objects and background. As $k$ increases, the semantics within a cluster diverge and form clusters with more precise semantics.
Figure A1: More qualitative results of the self-attention maps of DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
Figure A2: More qualitative results of zero-shot salient object segmentation. denotes the model with additional post-processing.
Figure A3: More qualitative results of semantic integrity of superpixel tokens with varying k. For each input image, the first row represents the embeddings of DeiT tokens, while the second row represents the embeddings of SuiT tokens.

### v3

@@
 A cornerstone of their success lies in the attention mechanism, which models relationships among tokens.
-While the tokenization process in NLP inherently ensures that a single token does not contain multiple semantics, the tokenization of Vision Transformer (ViT) utilizes tokens from uniformly partitioned square image patches, which may result in an arbitrary mixing of visual concepts in a token.
-In this work, we propose to substitute the grid-based tokenization in ViT with superpixel tokenization, which employs superpixels to generate a token that encapsulates a sole visual concept.
-Unfortunately, the diverse shapes, sizes, and locations of superpixels make integrating superpixels into ViT tokenization rather challenging.
-Our tokenization pipeline, comprised of pre-aggregate extraction and superpixel-aware aggregation, overcomes the challenges that arise in superpixel tokenization.
-Extensive experiments demonstrate that our approach, which exhibits strong compatibility with existing frameworks, enhances the accuracy and robustness of ViT on various downstream tasks.
-The source codes are available at: https://github.com/jangsoohyuk/SuiT
+While the tokenization process in NLP inherently ensures that each token maintains semantic integrity without containing multiple meanings, the grid-based tokenization of Vision Transformer (ViT) relies on uniformly partitioned square image patches, which may result in an arbitrary mixing of visual concepts within a token.
+In this work, we propose a novel tokenization pipeline that replaces the grid-based tokenization with superpixels, encouraging each token to capture a distinct visual concept.
+Unlike square image patches, superpixels are formed in varying shapes, sizes, and locations, making direct substitution challenging.
+To address this, our pipeline first generates pixel-level embeddings and efficiently aggregates them within superpixel clusters, producing superpixel tokens that seamlessly replace patch tokens in ViT.
+Extensive experiments demonstrate that our approach enhances the performance of ViT on various downstream tasks and introduces intriguing properties such as adaptive inference and semantic integrity in tokens.
 1 Introduction
@@
 Transformers [58], originally proposed for natural language processing (NLP), have demonstrated exceptional performance across diverse tasks such as machine translation [62, 22, 5], natural language understanding [30, 38], and text generation [46, 34].
-The transformative power of Transformers was again attested by the release of large language models [7], which continue to revolutionize NLP as we know it.
+The transformative power of Transformers was again attested by the release of large language models [7, 56].
 Similarly, the Vision Transformer (ViT) [14], which extended the Transformer architecture to image recognition, changed the landscape of computer vision research.
@@
 In NLP, the success of Transformers is commonly attributed to the attention mechanism, which captures contextual dependencies by modeling the relationship among input texts, which are expressed as tokens.
-Widely adopted tokenization methods in NLP, such as Byte Pair Encoding [46], WordPiece [47], and SentencePiece [29], first split each sentence into words, which are then broken down into more granular tokens.
-For instance, when tokenizing “Attention is All You Need,” each word is separated into: [“Attention,” “is,” “All,” “You,” “Need”].
-By definition, this initial word-level splitting of sentences prevents multiple words with different meanings from being fused into a single token.
-In contrast, fixed-size grid patches form the basis of tokenization in ViT and its variants [14, 50, 35].
+The widely adopted tokenization methods in NLP, such as Byte Pair Encoding [51], WordPiece [52] and SentencePiece [32], first divide each sentence into words based on whitespaces.
+For instance, when tokenizing “John loves Mary,” each word is separated into: [“John”, “ ”, “loves,” “ ”, “Mary”].
+This initial word-level splitting of sentences prevents multiple words with different meanings from being fused into a single token, preserving semantic integrity in tokens.
+In contrast, in computer vision, fixed-size grid patches form the basis of tokenization in ViT and its variants [14, 55, 39].
 As visualized in Figure 1, when naïvely sliced square image patches are used as tokens, a single token can simultaneously contain two or more visual concepts, e.g., eye and fur.
-Consequently, this approach in ViT diverges from the core principle of tokenization, which is to convert plain inputs into smaller yet more structured subunits for further processing.
-In this work, we employ the concept of superpixels to design a tokenization process that yields semantically preserving tokens that contain a single, isolated visual concept.
-A superpixel is a cluster of connected pixels in an image that share similar properties, such as color, texture, or position, and it is often used as a fundamental unit in image processing [2, 11, 33].
+This is analogous to splitting “John loves Mary” into arbitrary fixed-size fragments like [“Joh,” “n l,” “ove”, “s M”, “ary”], disregarding semantic boundaries.
+Consequently, this grid-based tokenization deviates from the semantic integrity inherent in NLP tokenization, leading to tokens with arbitrarily mixed visual concepts.
+In this work, we employ the concept of superpixels to design a tokenization process.
+A superpixel [2, 11, 36] is a cluster of connected pixels in an image that share similar properties, such as color, texture, or position, and is often used as a unit of semantic representation in various studies [17, 16].
+Building on this property, superpixels emerge as a viable alternative to square tokens for preserving semantic integrity.
 Figure 1 illustrates how leveraging superpixels alters tokenization of ViT.
-Grouping pixels into superpixels divides an image into cohesive regions that capture similar semantics, thereby preventing the mixture of unrelated semantic information in each token.
-Despite the advantage of superpixels in tokenization, the different shapes and sizes of each superpixel present significant challenges in utilizing them as ViT tokens [25, 59, 62].
-In ViTs, tokens are generated from fixed-size square patches at predefined locations through simple flattening and linear projection.
-Because superpixels consist of varying numbers of pixels, de-facto flattening and linear projection operations in ViT are incompatible with superpixels.
+Grouping pixels into superpixels divides an image into cohesive regions that capture similar semantics, thereby minimizing the mixture of unrelated semantic information in each token.
+Despite the potential advantages of superpixels in tokenization, the different shapes and sizes of each superpixel present significant challenges in utilizing them as a direct replacement for visual tokens [28, 67, 71].
+In ViTs, tokens are generated from fixed-size square patches at predefined locations through simple flattening and linear projection operations.
+Because superpixels consist of varying numbers of pixels, de facto flattening and linear projection operations in ViT are incompatible with superpixels.
 Moreover, the location of each superpixel changes dynamically depending on the image.
 The variability of superpixel locations makes it infeasible to directly apply the positional embeddings of ViT, which are designed for fixed locations.
-The inadequacy of original ViT operations necessitates novel techniques for addressing irregular geometry and composition to reap the inherent benefits of superpixels.
+The inadequacy of original ViT operations necessitates novel techniques to address irregular geometry and composition to reap the inherent benefits of superpixels.
 We thus propose a novel tokenization pipeline that overcomes the aforementioned problems and enables the effective incorporation of superpixels into ViT tokenization.
-The proposed pipeline consists of two major technical components: pre-aggregate feature extraction and superpixel-aware feature aggregation.
-The first stage prepares pre-aggregate features for the subsequent superpixel-aware aggregation, which utilizes pooling operations to remove the irregularity and variability of superpixels.
-Applying pooling operations directly on the input image may result in a severe loss of information.
-By training the pre-aggregate feature extractor and applying it prior to our feature aggregation, we effectively sidestep this potential information loss.
+The proposed pipeline consists of two major technical components: pixel-level embedding and superpixel-aware aggregation.
+The first stage generates pixel-level embeddings consisting of local and positional features.
+Second, these pixel-level embeddings are then aggregated according to superpixel clusters.
+The aggregation process employs pooling operations to remove the irregularity and variability of superpixels.
 Collectively, this two-stage process allows the proposed tokenization approach to leverage the benefits of superpixels while preserving important details of an image.
-We demonstrate the superiority of our Superpixel-Tokenized Vision Transformer (SuiT) compared to the baselines.
-SuiT outperforms in a variety of tasks including ImageNet-1K classification [12], segmentation [56], transfer learning [66], and self-supervised learning [9].
-We further analyze the properties that emerge using our tokenization pipeline by conducting extra examinations.
-Furthermore, we visualize that our proposed tokenization method indeed results in image tokens that preserve semantics more effectively than the grid-based tokenization method.
+We demonstrate the superiority of our Superpixel-Tokenized Vision Transformer (SuiT) compared to other baselines.
+SuiT outperforms in a variety of tasks including ImageNet-1K classification [12], transfer learning [75], and zero-shot segmentation [64].
+We further analyze intriguing properties that emerge using our tokenization pipeline, such as adaptive inference and semantic integrity.
 Our contribution can be summarized as follows:
-  • We propose an effective superpixel-based tokenization method that overcomes the challenges associated with the use of superpixels.
-  • We empirically show that our tokenization method produces tokens that better preserve semantic information than grid-based tokens.
-  • We demonstrate the effectiveness of our approach through experiments on image classification, transfer learning, self-supervised learning, and zero-shot segmentation.
+  • We propose an effective superpixel-based tokenization pipeline that overcomes the challenges associated with the use of superpixels.
+  • We empirically show that our tokenization pipeline induces notable properties including adaptive inference and semantic integrity in tokens.
+  • We demonstrate the effectiveness of our approach through experiments on image classification, transfer learning, and zero-shot segmentation.
 2 Related Work
@@
 Superpixels, which are perceptually meaningful clusters of pixels, have been a foundational concept in computer vision [2, 11, 36].
-Previous studies have attempted to use superpixels in several tasks such as segmentation [30, 37] and object detection [18], but these works did not demonstrate the broad applicability of superpixels across a range of vision tasks.
-Another recent work, STViT [23], applies superpixel-like clustering across all layers via supertoken attention, but operates at the token level, resulting in a coarser granularity than superpixels.
-CoC [36], borrows the concept of superpixels within neural networks but lacks further exploration of superpixel tokenization.
+Previous studies have attempted to use superpixels in several tasks such as segmentation [33, 41] and object detection [21].
+A recent work, STViT [26], applies superpixel-like clustering named supertoken attention in the Transformer backbone.
+However, unlike superpixels which are formed at the pixel level, clusters in STViT are formed at the square token level, resulting in coarser granularity.
+CoC [40] borrows the concept of superpixels within neural networks, but does not actually exploit superpixels in the network.
 Similarly, SPFormer [41] utilizes superpixel-based representations but requires specialized attention modules making it difficult to integrate with existing ViT architectures in a plug-and-play manner.
-In contrast to these approaches, our research explores a simple yet effective superpixel tokenization method that can be easily employed in a transformer architecture for a range of general-purpose tasks.
+These works all modify the transformer backbone to incorporate the concept of superpixels but do not explore their use in tokenization.
+In contrast, our work introduces a simple yet effective superpixel tokenization method that can be seamlessly integrated into the Transformer architecture without altering the backbone for a range of general-purpose tasks.
 Tokenization Methods
-While most ViT variants focus on improving backbone architectures and attention mechanisms using naive square-shaped patches, there have been several attempts in exploration of advanced tokenization strategies tailored for vision tasks.
-Quadformer [44] and MSViT [16] both propose adaptive tokenization strategies that dynamically adjust token resolutions based on image content.
-The studies SPiT [1] and sViT [28] are the most similar to our work in that they target only the tokenizing module while keeping the backbone intact and share similar research motivations.
-SPiT [1] focused on analyzing this tokenization from the explainable AI (XAI) perspective but did not explore its generalizability across various tasks.
-Also, SPiT [1] suffers from performance degradation without gradient features, as they lack of abundant features due to their learning free feature extraction.
-Moreover, its applicability to other learning paradigms, such as self-supervised learning [9], was not investigated.
-sViT [28] shared a similar motivation in aiming for semantically preserving tokenization, but it still divides tokens based on a naïve bounding box, failing to create truly semantically preserving tokens.
+While most ViT variants focus on improving backbone architectures and attention mechanisms using naïve square image patches, there have been several attempts in exploration of advanced tokenization strategies tailored for vision tasks.
+Quadformer [49] and MSViT [19] both propose adaptive tokenization strategies that dynamically adjust token resolutions based on the image content.
+The studies SPiT [1] and sViT [31] are the most similar to our work as they focus solely on the tokenization module while keeping the backbone intact and share similar research motivations.
+SPiT [1] uses superpixel tokenization but focuses on analyzing the tokenization from the perspective of explainable AI but do not demonstrate favorable performance compared to original ViT.
+sViT [31] shares a similar motivation in preserving semantic integrity during tokenization, but it still divides tokens based on a naïve bounding box, failing to create tokens with semantic integrity.
 Additionally, resizing these tokens into square shapes compromised the visual structure of objects within the tokens.
-These shortcomings in design leads leads to limited performance in various tasks.
-As described in the previous section, we propose a semantically preserving tokenization through superpixels in vision and investigate the performance of our method across various tasks and explore its adaptability to self-supervised frameworks.
+These shortcomings in design leads to limited performance in various tasks.
+As described in Section 1, we propose a tokenization that encourages semantic integrity in tokens through superpixels.
+We investigate the performance of our method across various tasks and explore its intriguing properties.
 Figure 2: Overview of our superpixel tokenization pipeline.
-Local features are extracted and combined with positional encodings, followed by superpixel-aware aggregation using mean and max pooling to produce superpixel tokens, which are fed into a Vision Transformer.
+Local features are extracted and combined with positional encodings, followed by superpixel-aware aggregation using average and max pooling to produce superpixel tokens, which are fed into Vision Transformer.
 3 Preliminaries
-In this section, we introduce the key concepts and notations that are integral to our work, which aims to construct semantically preserving ViT tokens with superpixels.
+In this section, we introduce the key concepts and notations that are integral to our work, which aims to construct ViT tokens that preserve semantic integrity using superpixels.
 In Section 3.1, we describe the mechanism of superpixel algorithms, and in Section 3.2, we go over the traditional ViT tokenization.
@@
 Superpixels are clusters of image pixels grouped together based on color, position, or other similarity measures.
-They provide a compact and meaningful representation of the image by grouping pixels into regions that align with its underlying structure.
+[2, 11, 36] They provide a compact and meaningful representation of the image by grouping pixels into regions that align with its underlying structure.
 Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, a superpixel algorithm $\mathcal{S}$ produces a superpixel index map $\mathcal{I}\in\mathbb{R}^{H\times W\times 1}$:
@@
 $\mathcal{C}_{k}=\{\textbf{x}^{(h,w)}\mid\mathcal{I}^{(h,w)}=k,0<k\leq K\},$ (2)
-where $\textbf{x}^{(h,w)}$ is the pixel at $(h,w)$. Each superpixel $\mathcal{C}_{k}$ consists of varying numbers of pixels with different shapes and sizes.
+where $\textbf{x}^{(h,w)}$ is the pixel at $(h,w)$. Each superpixel $\mathcal{C}_{k}$ consists of varying numbers of pixels with irregular spatial distributions.
 3.2 Tokenization in Vision Transformers
-In grid-based tokenization of conventional ViT and its variants [50, 35, 42], the input image $\mathbf{x}$ is divided into $N=\frac{H\times W}{P^{2}}$ non-overlapping patches of size $P\times P$, each represented as $\mathbf{x}^{(i)}\in\mathbb{R}^{P\times P\times C}$, where $\mathit{i}\in\{1,...,N\}$ corresponds to the index of each patch.
-Each patch $\mathbf{x}^{(i)}$ is flattened into a vector $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{P^{2}C}$ and linearly projected using an embedding matrix $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$, where $D$ is the dimension of tokens.
-In addition to this, positional embedding $\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$ of its location is injected to form the input token $\mathbf{z}^{(i)}=\mathbf{x}_{p}^{(i)}\mathbf{E}+\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$.
+In grid-based tokenization of conventional ViT and its variants [55, 39, 47], the input image $\mathbf{x}$ is divided into $N=\frac{H\times W}{P^{2}}$ non-overlapping patches of size $P\times P$, each represented as $\mathbf{p}^{(i)}\in\mathbb{R}^{P\times P\times C}$, where $\mathit{i}\in\{1,...,N\}$ corresponds to the index of each patch.
+Each patch $\mathbf{p}^{(i)}$ is flattened into a vector $\mathbf{p}_{flat}^{(i)}\in\mathbb{R}^{P^{2}C}$ and linearly projected using an embedding matrix $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$, where $D$ is the size of hidden dimension.
+In addition to this, positional embedding $\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$ of its location is injected to form the input token $\mathbf{z}^{(i)}=\mathbf{p}_{flat}^{(i)}\mathbf{E}+\mathbf{E}^{(i)}_{pos}\in\mathbb{R}^{D}$.
 The set of tokens $\mathbf{Z}=\{\mathbf{z}^{(i)}\mid i\in\{1,2,...,N\}\}$ which originates from patches as above, is given to the ViT models along with a learnable classification token $\mathbf{z}_{\text{cls}}$ for classification.
@@
 Flattening of superpixels with diverse shapes and varying numbers of pixels yields vectors of varying lengths, which hinder a simple linear projection.
-Second, unlike grid patches with fixed positions, superpixels are located at different positions depending on the image.
+Second, unlike grid patches with fixed positions, superpixels are located at different positions depending on the images.
 Consequently, the standard positional embedding, which relies on fixed positions, is inappropriate for superpixels.
 4 Method
-To address the aforementioned issues in Section 3, we propose a novel superpixel tokenization pipeline that effectively embeds each superpixel into a single token.
+To address the issues in Section 3, we propose a superpixel tokenization pipeline that embeds each superpixel into a single token.
 The overview of superpixel tokenization is illustrated in Figure 2.
-Our proposed pipeline incorporates two components: pre-aggregate feature extraction (Section 4.1) and superpixel-aware feature aggregation (Section 4.2).
-Our aggregation is based on pooling operations to address the inconsistent size and shape of superpixels.
+Our pipeline consists of two components: Pixel-level Embedding (Section 4.1) and Superpixel-aware Aggregation (Section 4.2).
+Our aggregation is based on pooling operations to address the varying sizes and shapes of superpixels.
 However, naïve application of such aggregation on the RGB space may lead to critical loss of information.
-To sidestep this potential risk, we train a pre-aggregate feature extractor and use its output as input for the superpixel-aware aggregation.
+(Table 4) To sidestep this potential risk, we use a pixel-level embedding layer to enrich pixel-level information before superpixel-aware aggregation.
 These two processes of the proposed tokenization method together exploit the benefits of superpixel-based tokenization while conserving critical image information.
-4.1 Pre-aggregate Feature Extraction
-Local Feature Extraction
+4.1 Pixel-level Embedding
+Local Features
 We start by extracting local features from the input image using a simple convolutional block.
@@
 Positional features play a crucial role in Transformers, as they are unaware of the order of input tokens.
-While the original Transformer proposed to use sinusoidal positional encodings with pre-defined frequencies [51], ViTs opted to use learnable positional embeddings for each pre-defined locations [14, 50].
+While the original Transformer proposed to use sinusoidal positional encodings (PE) with pre-defined frequencies [58], ViTs opted to use learnable positional embeddings for each pre-defined locations [14, 55].
 Superpixels have complex positional information of high granularity.
-Using the learnable positional embeddings as in ViT is inadequate, as it would require excessive number of parameters, i.e., $H\times W$ learnable embeddings each of $D$-dimensions.
+Using the learnable positional embeddings along every pixels as in ViT is inadequate, as it would require excessive number of parameters, i.e., $H\times W$ learnable embeddings each of $D$-dimensions.
 To keep it parameter-efficient, we adopt the sinusoidal positional encoding with learnable frequencies, which is known to capture high frequency information efficiently [54].
@@
 Projection
-Given the local features $\mathbf{F}_{local}$ and the positional features $\mathbf{F}_{pos}$, we integrate the two to obtain our pre-aggregate features.
-The two feature maps $\mathbf{F}_{local}$ and $\mathbf{F}_{pos}$ are concatenated and passed to a linear projection layer $\mathcal{P}$:
-$\mathbf{F}=\mathcal{P}(\langle\mathbf{F}_{local},\mathbf{F}_{pos}\rangle),$ (4)
+Given the local features $\mathbf{F}_{local}$ and the positional features $\mathbf{F}_{pos}$, we integrate the two to obtain our pixel-level embeddings.
+The two feature maps $\mathbf{F}_{local}$ and $\mathbf{F}_{pos}$ are concatenated and passed through a linear projection layer, represented as a weight matrix $\mathbf{W}_{P}$:
+$\mathbf{F}=\langle\mathbf{F}_{local},\mathbf{F}_{pos}\rangle\mathbf{W}_{P}$ (4)
 […]
-4.2 Superpixel-aware Feature Aggregation
-To aggregate the extracted pre-aggregate features within each superpixel into a representative embedding, we employ average and max pooling, which effectively handles features of varying numbers.
-The pre-computed superpixel index map $\mathcal{I}$ of an input image x is used as the guideline for determining which features belong to which superpixels.
-Given the $D/2$-dimensional pre-aggregate features, we perform average pooling and max pooling within each superpixel cluster.
-This results in two feature vectors per superpixel: average features $\mathbf{z}_{\text{avg}}$ and max features $\mathbf{z}_{\text{max}}$, each of $D/2$-dimensions:
+4.2 Superpixel-aware Aggregation
+To form an embedding of each superpixel, we aggregate the pixel-level embeddings within each superpixel.
+The superpixel index map $\mathcal{I}$, acquired by the superpixel algorithm, serves as a guideline in determining which embeddings correspond to which superpixels.
+Based on $\mathcal{I}$, we aggregate $D/2$-dimensional pixel-level embeddings within each superpixel via average and max pooling, which effectively handles embeddings of varying numbers.
+This results in two embedding vectors per superpixel: average embeddings $\mathbf{z}_{\text{avg}}$ and max embeddings $\mathbf{z}_{\text{max}}$, each of $D/2$-dimensions:
 $\begin{split}\mathbf{z}^{(k)}_{\text{avg}}&=\frac{1}{|\mathcal{C}_{k}|}\sum_{\textbf{x}^{(h,w)}\in\mathcal{C}_{k}}\mathbf{F}^{(h,w)},\\ \mathbf{z}^{(k)}_{\text{max}}&=\max_{\textbf{x}^{(h,w)}\in\mathcal{C}_{k}}\mathbf{F}^{(h,w)},\end{split}$ (5)
-where $|\mathcal{C}_{k}|$ and $\mathbf{F}^{(h,w)}$ denote the number of pixels within superpixel $\mathcal{C}_{k}$ and the pre-aggregate feature at coordinate $(h,w)$, respectively.
-Average pooling and max pooling extract complementary features from each superpixel.
-Average pooling captures the overall characteristics common in the superpixel, while max pooling identifies the most salient features that may be obscured by averaging alone.
+where $|\mathcal{C}_{k}|$ and $\mathbf{F}^{(h,w)}$ denote the number of pixels within superpixel $\mathcal{C}_{k}$ and the pixel-level embeddings at coordinate $(h,w)$, respectively.
+Average pooling and max pooling extract complementary embeddings from each superpixel.
+(5) Average pooling captures the overall characteristics common in the superpixel, while max pooling identifies the most salient features that may be obscured by averaging alone.
 We concatenate $D/2$-dimensional vectors from each pooling method along the channel dimension, yielding a $D$-dimensional token $\mathbf{z}^{(k)}$:
@@
 5 Experiments
-Size Model # Params.
-GMACs # tokens Top-1 Tiny DeiT 5.7M 1.26 196 72.2 SuiT 5.5M 0.79 100 72.2 1.26 171 75.3 1.44 196 75.7 Small DeiT 22.1M 4.61 196 79.8 SuiT 21.7M 3.64 132 79.8 4.60 172 80.5 5.20 196 80.9 Base DeiT 86.8M 17.6 196 81.8 SuiT 86.0M 16.4 160 81.8 17.6 173 82.0 19.7 196 82.1
-Table 1: Comparison of classification performance on ImageNet-1K against DeiT.
-The performance of SuiT is reported with varying token counts.
-Underlined values indicate those that match DeiT’s GMACs or token counts.
+Size Model Weight-Init. # Params. GMACs Top-1 (%)
+Tiny DeiT [55] Random 5.7M 1.26 72.2
+SPiT [1] - - -
+SuiT 5.5M 1.44 75.7
+MSViT [19] IN-21K - 0.98 75.5
+Quadformer [49] - - -
+SuiT 5.5M 1.44 76.1
+Small DeiT [55] Random 22.1M 4.61 79.8
+SPiT [1] - - 75.0
+SuiT 21.7M 5.20 80.9
+MSViT [19] IN-21K - 4.43 82.0
+Quadformer [49] - 4.71 80.8
+SuiT 21.7M 5.20 82.6
+Base DeiT [55] Random 86.8M 17.6 81.8
+SPiT [1] - - 80.4
+SuiT 86.0M 19.7 82.3
+MSViT [19] IN-21K - - -
+Quadformer [49] - 17.7 84.4
+SuiT 86.0M 19.7 84.5
+Table 1: Comparison of classification performance on ImageNet-1K [12].
+Following prior studies, we condcuted experiments on two different settings for fair comparisons.
+In the “Weight-Init.” column, “Random” refers to from-scratch training on ImageNet-1K [12], while “IN-21K” refers to training from ViT weights pre-trained on ImageNet-21K [48].
+denotes that the method was initialized with IN-21K pre-trained + IN-1K fine-tuned ViT weights.
+Cells marked ‘-’ means that the value is not available.
+Figure 3: Adaptive inference.
+Under the same image resolution (=224), DeiT [55] is forced to use a fixed # tokens (=196), whereas both SuiT and SPiT [1] can adaptively adjust the number of tokens during inference.
+Notably, across all token counts and model scales, SuiT consistently achieves the best performance compared to baselines under the same computational cost.
 5.1 Experimental Settings
-We validate our tokenization method by applying it to ViTs, dubbed as SuiT.
-We experiment on three model scales, Tiny ($D=192$), Small ($D=384$), and Base ($D=768$).
-The values $d_{local}$ and $d_{pos}$ are both set to $D/4$.
-We employ FastSLIC [4] as our superpixel algorithm due to its high speed.
-We use a stride of 2 in the convolutional layer of local feature extraction, which produces feature maps of 2×2$\times$2 × smaller size than the input.
-The positional feature map $\mathbf{F}_{pos}$ and superpixel index map $\mathcal{I}$ are scaled to the same size accordingly.
-This design choice was based on the observation that using a stride of 1 offered no notable performance improvements while significantly increasing computational costs.
-Further details are described in Appendix.
-5.2 Comparison to Baseline
-As our approach focuses on tokenization without modifying the backbone network, we focus on comparison with our baseline, DeiT [50] on ImageNet-1K classification [12].
-We mostly follow the experimental settings from DeiT, with minimal changes for fair comparison.
-Table 1 compares the classification performance of SuiT to that of DeiT.
-For each model size, the table provides a detailed comparison across key aspects: classification accuracy, computational cost measured in GMACs, and the number of tokens processed.
-This adaptive inference capability arises from the flexibility to adjust the number of superpixel tokens, enabling users to customize the configuration based on specific needs.
-When achieving comparable performance, SuiT exhibits substantially higher efficiency than DeiT.
-At equivalent GMACs, SuiT delivers superior accuracy.
-Maintaining an identical number of tokens further boosts its accuracy,
-Tiny Small Base
-Quadformer [44] - 80.8 84.4
-MSViT [16] 75.5 82.0 -
-SuiT (Fine-tuned) 76.1 82.6 84.5
-SPiT [1] - 75.0 80.4
-SuiT (Scratch) 75.7 80.9 82.1
-Table 2: Comparison with existing studies on tokenization.
-Experiment above reports the Top-1 accuracy on ImageNet-1K.
-The numbers for Quadformer-Tiny, MSViT-Base and SPiT-tiny are intentionally left blank, as they are unavailable.
+We validate our tokenization method by applying it to ViT, referring to the resulting model as SuiT.
+As baselines, we include DeiT [55] and three tokenization-focused approaches—SPiT[1], Quadformer[49], and MSViT[19]—while keeping the backbone unchanged.
+We evaluate three model scales: Tiny $(D=192)$, Small $(D=384)$, and Base $(D=768)$, setting the local and positional feature dimensions to $D/4$.
+All models process $224\times 224$ images during training and evaluation.
+For pixel-level embedding, we use a stride-2 convolution, reducing feature map size by half.
+The positional feature map $\mathbf{F}_{pos}$ and superpixel index map $\mathcal{I}$ are down-scaled accordingly.
+A preliminary study showed that using stride 1 yielded negligible gains at a much higher computational cost, so we opted for stride 2.
+Among the baseline methods, SPiT [1] is the most closely related to ours, as it also employs superpixels in tokenization.
+Meanwhile, Quadformer [49] and MSViT [19] do not utilize superpixels but improve the tokenization process through mixed-scale tokenization.
+Following the settings of each method, we conduct experiments in two different settings.
+For DeiT [55] and SPiT [1], we train and evaluate the models using the standard ImageNet-1K (IN-1K) [12] classification setting [55, 20, 39, 63, 68], where all models are trained from scratch.
+In contrast, for Quadformer [49] and MSViT [19], we follow their experimental setups.
+Specifically, MSViT initializes the transformer backbone with a ViT model pre-trained on ImageNet-21K (IN-21K) [48], while Quadformer uses a ViT model that is first pre-trained on IN-21K and then further fine-tuned on IN-1K.
+In our experiments, we follow the setting of Quadformer.
+For the superpixel algorithm, we employ SLIC [2], but in practice, we use its faster implementation, FastSLIC [4] due to its computational efficiency.
+We configure SLIC [2] with its default setting, a compactness of 10, a maximum of 10 iterations, and 196 superpixels.
+Note that the value 196 is simply the maximum number of superpixels, and the actual number of superpixels varies across images.
+For example, on the ImageNet-1K [12] validation set, the average number of superpixels is approximately 191.9 ± 6.0.
+5.2 Comparison to Baselines
+Table 1 compares the classification performance of SuiT with baselines in two different settings.
+For each model size, the table provides a detailed comparison across key aspects: number of total parameters of the model (# Params.), computational cost measured in GMACs (GMACs), and the classification accuracy (Top-1).
+The results in Table 1 demonstrate that SuiT consistently outperforms the baselines in both experimental settings.
+Specifically, when trained from scratch with randomly initialized weights, SuiT achieves performance gains across all model sizes compared to DeiT [55]: Tiny (+3.5%p), Small (+1.1%p), and Base (+0.5%p).
+Additionally, SuiT surpasses SPiT [1], the most closely related superpixel tokenization baseline, by substantial margins: Small (+5.9%p) and Base (+1.9%p).
+In the setting where weights are initialized with IN-21K [48], SuiT consistently outperforms Quadformer [49] and MSViT [19] across all model scales.
+Notably, in the Tiny model, SuiT with random initialization achieves superior performance compared to MSViT [19] initialized with IN-21K [48] pre-trained weights.
+Furthermore, in the Small model, our approach trained with random initialization even surpasses Quadformer [49] initialized with IN-21K [48] + IN-1K [12] pre-trained weights.
+5.3 Adaptive Inference
+Figure 3 highlights the adaptive inference capability of SuiT.
+This figure illustrates the performance variation with respect to the number of tokens during inference, where SPiT [1] is plotted based on the average token count, while SuiT is plotted using the maximum token count.
+Unlike DeiT [55], which relies on a fixed number of tokens, SuiT can dynamically adjust the number of superpixel tokens during inference, enabling more efficient computation while maintaining strong performance.
+Notably, SuiT outperforms DeiT [55] even with fewer tokens, requiring only 100 tokens compared to DeiT’s 196 to achieve the same performance in the tiny model setting.
+This highlights its ability to extract meaningful visual representations with higher efficiency.
+Additionally, SuiT can further increase the number of superpixels at inference time, effectively scaling accuracy as needed.
+Notably, despite being trained with 196 superpixels, increasing to 256 at inference improves accuracy across different model scales: from 75.7% to 75.9% in the Tiny model, 80.9% to 81.2% in the Small model, and 82.3% to 82.5% in the Base model.
+This contrasts with SPiT [1], which, despite its ability to adjust token counts, fails to deliver competitive performance across all token settings.
+In specific, SPiT-B [1] achieves only 64.0% accuracy with an average of 46 tokens, while increasing the token count to an average of 774 marginally improves performance to 79.2%.
+Note that SPiT [1] does not allow direct control over the number of tokens but instead relies on a “level” hyperparameter, limiting fine-grained adjustments and restricting evaluations to extreme settings (46 and 774 tokens).
+These findings underscore the strength of SuiT’s adaptive inference, allowing it to flexibly balance computational cost and accuracy, making it a more scalable and efficient alternative to both DeiT [55] and SPiT [1].
+5.4 Transfer Learning
+We verify the generalization ability of SuiT when transferred to various downstream tasks.
+We follow the standard fine-tuning protocol and default settings of DeiT [55] for fair comparison, by fine-tuning the ImageNet-1K pre-trained SuiT to each downstream task specifically.
+We test our models on iNaturalist [57], Flowers102 [44], and StanfordCars [10].
+Further details can be found in the Appendix A2.2.
 Size Model $\text{INat}_{18}$ $\text{INat}_{19}$ Cars Flowers
@@
 SuiT 74.0 78.4 92.1 96.9
-DINO-ViT 72.6 78.6 93.0 97.1
-DINO-SuiT 72.8 79.4 93.8 97.2
 Table 2: Results of transfer learning different datasets in terms of top-1 accuracy results.
 SuiT outperforms ViT in most downstream tasks in both supervised and self-supervised settings.
-All results for DeiT and DINO are reproduced following the training recipes provided by the authors.
 : The reported Flowers102 performance could not be reproduced in the exact setting, so we compare with our reproduced results.
@@
 denotes that the model is trained at a different training setting to others, i.e., with different resolution ($256$ vs $224$), data augmentation techniques and learning method (supervised vs self-supervised).
-5.3 Comparison to Existing Tokenization Methods
-We evaluate our tokenization method against existing approaches under two scenarios used in existing tokenization studies.
-For a fair comparison, we follow the setup of  [44, 16], which fine-tunes a ViT pretrained on ImageNet-21K [43].
-Additionally, we include SPiT  [1], a concurrent work leveraging superpixels for tokenization, which trains its model from scratch on ImageNet-1K.
-Both experiments are conducted on ImageNet-1K classification tasks.
-As reported in Table 2, our method consistently outperforms existing tokenization methods across all scales of model size, showing its efficacy.
-On fine-tuning experiments following [44, 16], SuiT outperforms existing tokenization methods for fine-grained tokens, such as Quadformer [44] and MSViT [16].
-In experiments for training from scratch following [1], SuiT consistently outperforms SPiT, another work on superpixel-based tokenization, by a noticeable margin.
-These results verify the excellence of our tokenization pipeline compared to prior work.
-5.4 Transfer Learning
-We verify the generalization ability of SuiT when transferred to various downstream tasks.
-We follow the standard fine-tuning protocol and default settings of DeiT [50] for fair comparison, by fine-tuning ImageNet-1K pre-trained SuiT to each downstream task specifically.
-We test our models on iNaturalist [22], Flowers102 [39], and StanfordCars [10].
-Further details can be found in the Appendix.
-Table 3 presents a comparison of the classification accuracy between DeiT and SuiT when fine-tuned on various downstream tasks.
+Table 2 presents a comparison on the classification accuracy between DeiT and SuiT when fine-tuned on various downstream tasks.
 Across various tasks and model scales, SuiT consistently achieves higher accuracy than DeiT.
-Notably, on iNaturalist2018, SuiT-Small outperforms DeiT-Small by 0.8%p, demonstrating its superior ability to effectively transfer learned features, particularly when ample data is available.
-Figure 3: Self-attention map comparison between DINO-ViT and DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
+Notably, iNaturalist2018 [25], a dataset of sufficient scale, SuiT-Small outperforms DeiT-Small by 0.8%p, highlighting its superior ability to transfer learned features effectively especially in data-rich scenarios.
+These results indicate that SuiT shows improved generalizability than DeiT.
+This positions SuiT a strong choice for transfer learning across diverse domains.
+The consistently superior performance of SuiT, regardless of model size or dataset complexity, further underscores its suitability for real-world applications.
+5.5 Zero-shot Salient Object Segmentation
+We conduct zero-shot salient object segmentation to demonstrate the advantages of superpixel tokens in dense prediction tasks.
+We adopt DINO [9] models for this experiment.
+We evaluate the performance on ECSSD [65], DUTS [61] and DUT-OMRON [66] using the TokenCut framework [64].
+Superpixel tokenization significantly boosts DINO-ViT’s performance across most metrics, even without post-processing, as shown in Table 3.
+Unlike DINO-ViT, SuiT maintains minimal performance differences between processed and unprocessed results.
+Compared to SPiT [1], SuiT achieves competitive performance, though SPiT was trained on higher-resolution images.
+Moreover, as Figure 4 shows, DINO-SuiT detects all salient objects, even in multi-object scenarios, whereas DINO-ViT struggles with rectangular masks and incomplete detections.
+The visualization of self-attention maps of DINO-SuiT can be found in the Appendix A2.3.
 Figure 4: Qualitative results of zero-shot salient object segmentation.
@@
 DINO-SuiT successfully detects salient objects both in single- and multi-object scenarios without any post-processing.
-These results indicate that SuiT shows improved generalizability than DeiT.
-This positions SuiT a strong choice for transfer learning across diverse domains.
-The consistently superior performance of SuiT, regardless of model size or dataset complexity, further underscores its suitability for real-world applications.
-5.5 Self-supervised Learning
-We further study the applicability of SuiT in self-supervised scenario.
-Specifically, we train SuiT on ImageNet-1K with DINO [9], a well-known self-supervised learning approach, chosen for its proven effectiveness and popularity.
-5.5.1 Transfer Learning
-After pre-training SuiT with DINO, we evaluate its performance when transferred to downstream tasks to confirm the quality of learned features.
-As shown in the lower section of Table 3, DINO-SuiT consistently surpasses DINO-ViT across all downstream tasks.
-This demonstrates that SuiT effectively supports self-supervised scenarios, further establishing its strength as a versatile pipeline for a wide range of vision tasks.
-5.5.2 Zero-shot Salient Object Segmentation
-It is well known that the attention map of ViTs trained with DINO effectively attends to object regions without explicit supervision [9].
-As shown in Figure 3, while DINO-ViT’s self-attention maps are limited to square patch-level details, those of DINO-SuiT even capture finer structures of salient objects.
-To demonstrate that this fine-grained property synergistically enhances the salient object segmentation capabilities of DINO [9], we measure the segmentation performance on three benchmark datasets, ECSSD [57], DUTS [54] and DUT-OMRON [58] following the works of Wang et al.
-[56], Aasan et al.
-[1].
-We conduct our experiments on the TokenCut framework [56], which is a normalized graph cut algorithm that extracts saliency mask from attention map.
-Superpixel tokenization significantly improves DINO-ViT [9] performance across most metrics without post-processing, achieving results comparable to DINO-ViT with post-processing.
-Notably, as shown in Table 4, our model exhibits minimal performance differences between post-processed and non-post-processed cases, unlike DINO-ViT.
-Compared to SPiT [1], SuiT demonstrates competitive performance across three datasets.
-However, SPiT [1] does not show its applicability to self-supervised methods such as DINO and is trained with larger resolution images.
-Qualitative comparisons in Figure 4 highlight that DINO-SuiT detects all salient objects, even in multi-object scenarios, whereas DINO-ViT shows limitations such as rectangular masks without post-processing and incomplete object detection.
+5.6 Semantic Integrity of Superpixel Tokens
+To assess the semantic preservation of our superpixel tokens, we perform K-means clustering [29] in the token embedding space and compare the results with those of patch tokens.
+As shown in Figure 5, our superpixel tokens cluster based on semantic similarity, whereas DeiT’s patch tokens lack consistency.
+For example, in the first row, tokens from duckling regions are grouped together, while DeiT’s patch tokens fail to capture clear semantic meaning.
+This suggests that superpixel tokens better preserve semantic integrity.
+SPiT, which also utilizes superpixels in tokenization, exhibits similar behavior to SuiT.
+However, SPiT clusters the shadows together with the ducklings in the first row, whereas those of SuiT selectively groups only the ducklings.
+This further supports that our tokenization method preserves stronger semantic coherence.
+Additional results can be found in the Appendix A3.
 Figure 5: Comparison of K-means clustering results for patch tokens from DeiT and superpixel tokens from SuiT, both supervisedly trained on ImageNet-1K, with K=6.
 Superpixel tokens in SuiT tend to cluster by semantic meaning, while ViT patch tokens mostly cluster by positional information.
-6 Analysis
-6.1 Semantic Integrity of Superpixel Tokens
-To verify the improved semantic preservation of our superpixel tokens, we perform K-means clustering [26] on tokens embedding space.
-We compare the clustering results of superpixel tokens with grid-based tokens.
-As shown in Figure 5, our superpixel tokens are grouped based on semantic similarity rather than spatial proximity, unlike the patch tokens from DeiT, which lack consistency within clusters.
-For example, in the first row, tokens corresponding to the regions of ducklings are grouped together, whereas clusters from DeiT’s patch tokens fail to convey clear meaning.
-This analysis suggests that superpixel tokens are better at preserving semantics.
-Additional results are provided in the Appendix.
-Figure 6: Class-identifiability analysis results of DeiT and SuiT. We show the proportion of class-identifiable [CLS] tokens and image tokens for SuiT and DeiT across all layers.
+Local feats. PE Freqs. PE Injection Top-1
+random init. concat+proj 75.7
+random init. concat+proj 66.5
+No PE 71.5
+pre-defined concat+proj 73.0
+random init. concat 73.8
+random init. addition 75.2
+Table 4: Ablation study on technical components of pre-aggregate feature extraction.
+Avg Std Softmax Max Top-1
+75.7
+73.2
+75.5
+Fail
+75.5
+Table 5: Ablation on our aggregation method. ‘Fail’ indicates that the corresponding approach was unsuccessful in training.
+Compactness Iters Algorithm Top-1
+10 10 Fast-SLIC 75.7
+1 10 Fast-SLIC 74.4
+100 10 Fast-SLIC 75.5
+Square N/A N/A 75.1
+10 1 Fast-SLIC 75.4
+10 20 Fast-SLIC 75.6
+0.001 N/A Watershed 73.9
+Table 6: Ablation on the superpixel algorithm.
+5.7 Ablation Studies
+Pixel-level Embeddings
+The following ablation studies on our proposed components are conducted using SuiT-Tiny.
+Table 4 highlights the importance of components in obtaining pixel-level embeddings.
+First, using raw 3-channel RGB features (row 2) leads to a significant performance drop, confirming the need for higher-dimensional local features.
+Second, removing positional information (row 3) reduces performance but retains partial class prediction through local cues.
+Experiments also reveal that pre-defined positional frequencies (row 4) under-perform compared to learned frequencies, emphasizing the need for vision-specific customization.
+Lastly, our results show that the combination of concatenation and projection more effectively integrate local and positional features than simple concatenation (row 5) or addition (row 6) used in prior studies [58, 14, 55].
+Aggregation
+Table 5 presents the analysis of our aggregation method, which employs cardinality-independent operations.
+Removing max pooling (row 2) results in a notable performance drop, and removing average pooling (row 3) causes a slight decrease.
+We also tested standard deviation as an alternative to max pooling, but it proved numerically unstable and unsuitable for training (row 4).
+Softmax pooling (row 5), which computes a weighted sum of values within each superpixel, showed promising results but slightly underperformed compared to max pooling.
+This may be due to the softmax operation suppressing prominent features crucial for recognition, leading to less discriminative representations.
+Superpixel Method
+In Table 6, we present a study regarding the superpixel methodology.
+First, we examine the compactness parameter in SLIC, which controls the trade-off between spatial proximity and pixel content in superpixel formation (rows 2-4).
+Higher values produce more regular shapes but reduce semantic relevance.
+We compare compactness values of 1 (row 2) and 100 (row 3), along with square-shaped tokens (row 4) as an upper-bound of high compactness.
+Keeping other components unchanged, we find that increasing compactness degrades performance, supporting the importance of semantic integrity.
+However, an extremely low value (e.g., 1) also harms performance, likely due to insufficient spatial coherence.
+Our framework relies on SLIC, which iteratively clusters pixels.
+We try different number of iterations in clustering as shown in Table 6, but the number of iterations had little impact on performance (rows 5 and 6).
+We also test Watershed [43] (row 7), which results in a slight performance drop but remains robust.
+6 Limitation and Future Work
+Our method, in its current form, is limited to the vanilla ViT architecture[14].
+Yet, since the ViT architecture is the most prevalent choice for vision architecture in many applications, e.g., CLIP [47] and LLaVA [37] in multi-modal learning, we believe that building an improvement for the ViT architecture is valuable.
+Applying superpixel tokenizations to other ViT variants, e.g., Swin Transformer [39] remains a topic for future study.
+7 Conclusion
+In this work, we proposed a novel superpixel-based tokenization method for ViTs as an attempt to improve the visual tokenization process.
+By leveraging superpixels, the proposed tokenization obtains tokens preserving semantic integrity, overcoming the issue of mixed semantics in grid-based patches.
+To incorporate superpixels that are irregular in shape, size, and location, we proposed a two-stage pipeline consisting of pixel-level embedding and superpixel-aware aggregation.
+Extensive experiments demonstrated the effectiveness and intriguing properties of our method.
+We hope our promising results inspire further exploration of token design in ViTs, fostering advancements in both model performance and interpretability.
+Supplementary Material
+Figure A1: Class-identifiability analysis results of DeiT and SuiT. We show the proportion of class-identifiable [CLS] tokens and image tokens for SuiT and DeiT across all layers.
 Size Model ImageNet-A ImageNet-O
@@
 SuiT 29.5 26.7
-Table 5: Experimental results of SuiT and DeiT on ImageNet-A and ImageNet-O. We report accuracy on ImageNet-A and AUPR on ImageNet-O, following  [21].
-6.2 Class Identifiability Analysis
-Vilas et al.
+Table A1: Experimental results of SuiT and DeiT on ImageNet-A and ImageNet-O. We report accuracy on ImageNet-A and AUPR on ImageNet-O, following  [24].
+A1 Additional Experiments
+A1.1 Class Identifiability Analysis
 [59] examined class-specific encoding in supervised ViTs by projecting intermediate representations through the trained classifier head into a class-specific embedding space, similar to the logit lens method in NLP [45, 6].
-Building on this, we evaluate class-specific encoding across SuiT’s layers by projecting token representations through the classifier head and measuring the proportion of tokens correctly predicting the target class.
-Figure 6 shows that SuiT’s classification token shows higher class identifiability in the earlier layers compared to DeiT.
+Building on this, we evaluate class-specific encoding across SuiT’s layers by projecting token representations through the classifier head and measuring the proportion of tokens correctly predicting the target class on IN-S-919 [18].
+Figure A1 shows that SuiT’s classification token shows higher class identifiability in the earlier layers compared to DeiT.
 This can be attributed to the semantic preserving nature of superpixel tokenization, which aids in aggregating class-specific encoding in early layers.
@@
 This divergence alludes that SuiT and DeiT adopt distinctive strategies to aggregate class-specific features at different network depths.
-6.3 Enhanced Robustness
+A1.2 Enhanced Robustness
 We examine the robustness and out-of-distribution (OOD) generalization ability of SuiT in ImageNet variants, ImageNet-A [13] and ImageNet-O [24].
 ImageNet-A is a subset of ImageNet-1K curated to mislead classifiers and test model robustness, while ImageNet-O, composed of OOD samples not in ImageNet-1K, evaluates the OOD detection capabilities of models.
-In Table 5, SuiT consistently outperforms DeiT on both datasets across model variants.
+In Table A1, SuiT consistently outperforms DeiT on both datasets across model variants.
 These findings indicate that our superpixel tokenization improves model robustness in challenging scenarios by mitigating the model’s reliance on features that are easily learned but lack generalizability.
-6.4 Ablation Studies
-Local feats. PE Freqs. PE Injection Top-1
-random init. concat+proj 75.7
-random init. concat+proj 66.5
-No PE 71.5
-predefined concat+proj 73.0
-random init. addition 75.2
-random init. concat 73.8
-Table 6: Ablation study on technical components of pre-aggregate feature extraction.
-Avg Std Softmax Max Top-1
-75.7
-73.2
-75.5
-Fail
-75.5
-Table 7: Ablation on our aggregation method. ‘Fail’ indicates that the corresponding approach was unsuccessful in training.
-Pre-aggregate Features
-The following ablation studies on our proposed components are conducted using SuiT-Tiny.
-Table 6 highlights the importance of components in pre-aggregate feature extraction.
-First, using raw 3-channel RGB features leads to a significant performance drop, confirming the need for higher-dimensional local features.
-Second, removing positional information reduces performance but retains partial class prediction through local cues.
-Experiments also reveal that predefined positional frequencies under-perform compared to learned frequencies, emphasizing the need for vision-specific customization.
-Lastly, our results show that the combination of concatenation and projection more effectively integrate local and positional features than simple concatenation or addition used in prior studies [51, 14, 50].
-Aggregation
-Table 7 presents the analysis of our aggregation method, which employs cardinality-independent operations.
-Removing max pooling results in a notable performance drop, and removing average pooling causes a slight decrease.
-We also tested standard deviation as an alternative to max pooling, but it proved numerically unstable and unsuitable for training.
-Softmax pooling, which computes a weighted sum of values within each superpixel, showed promising results but slightly underperformed compared to max pooling.
-This may be due to the softmax operation suppressing prominent features crucial for recognition, leading to less discriminative representations.
-7 Conclusion
-In this work, we proposed a novel superpixel-based tokenization method for ViTs as an attempt to improve the visual tokenization process.
-By leveraging superpixels, the proposed tokenization obtains semantic-preserving tokens, overcoming the issue of mixed semantics in grid-based patches.
-To incorporate superpixels that are irregular in shape, size, and location, we proposed a two-stage pipeline consisting of pre-aggregate feature extraction and superpixel-aware feature aggregation.
-Extensive experiments demonstrated the effectiveness of our method on diverse tasks, e.g., supervised classification, transfer learning, and self-supervised learning.
-Additional analytical findings provided meaningful insights into the mechanism and emergent properties of superpixel tokenization.
-We hope our promising results inspire further exploration of token design in ViTs, fostering advancements in both model performance and interpretability.
-Supplementary Material
-A1 Implementation Details
-A1.1 ImageNet-1K classification
-A1.1.1 Training from scratch
+A2 Implementation Details
+A2.1 ImageNet-1K classification
+A2.1.1 Training from scratch
 The settings for experiments from scratch is presented in Table A3.
@@
 We thus opt to use a stronger regularization term.
-Similar observations were made in Transfer Learning, which is also discussed in the Section  A1.2.
-A1.1.2 Fine-tuning from Pre-trained weights
+Similar observations were made in Transfer Learning, which is also discussed in the Section  A2.2.
+A2.1.2 Fine-tuning from Pre-trained weights
 […]
-A1.2 Transfer Learning
+Figure A2: Self-attention map comparison between DINO-ViT and DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.
+A2.2 Transfer Learning
 Here, we list the training settings used for transfer learning.
@@
 Cutmix prob. 1.0
-Erasing prob. 0.1
+Erasing prob. 0.25
 Table A2: Settings of SuiT on ImageNet-1K classification, training from scratch. We mostly follow the settings open to public by the authors [55].
@@
 Cutmix prob. 1.0
-Erasing prob. 0.1
+Erasing prob. 0.25
 Table A3: Settings of SuiT on ImageNet-1K classification, fine-tuning from pre-trained weights. We mostly follow the settings open to public by the authors [55].
@@
 Table A5: Training settings for transfer learning DINO pre-trained SuiT. We mostly follow the settings open to public by the authors [9].
-A1.3 Self-Supervised Learning
+A2.3 Self-Supervised Learning
 For DINO [9] pre-training, we follow the exact training hyper-parameters used to pre-train ViTs, without any tuning, extra training techniques or augmentations.
 We train DINO-SuiT for 300 epochs, and transfer the pre-trained model to downstream tasks following the settings mentioned in Section A2.2.
-Additional self-attention maps of DINO-SuiT can be found in Figure A1.
-A1.4 Zero-shot Salient Object Segmentation
+Additional self-attention maps of DINO-SuiT can be found in Figure A2 and A4.
+Figure A3: Visulization of superpixel tokenization using various parameters.
+A2.4 Zero-shot Salient Object Segmentation
 In the TokenCut [64] algorithm, we searched for and used the best $\tau$ value for both DINO-ViT and DINO-SuiT.
@@
 More qualitative results are shown in Figure A5.
-A2 Semantic Integrity of Superpixel Tokens
+A3 Semantic Integrity of Superpixel Tokens
 In our semantic integrity analysis, we performed $k$-means clustering on each token embeddings from DeiT [55] and SuiT.
@@
 As $k$ increases, the semantics within a cluster diverge and form clusters with more precise semantics.
+A4 Visualization of Superpixel Tokenization
+We visualize how an image can be tokenized using various parameters in superpixel generation process at Figure A3.
+We show how the tokens change when setting different number of superpixels to be generated (compactness set to 10) and when setting different number of compactness (number of superpixels set to 196).
 Figure A4: More qualitative results of the self-attention maps of DINO-SuiT. All maps are visualized by averaging the [CLS] token’s self-attention from the final layer across all heads.